Can I get it solved step by step please and not in excel.

Thank you

You purchased shares of a mutual fund at a price of $15 per share at the beginning of the year and paid a front-end load of 5.75%. If the securities in which the fund invested increased in value by 12% during the year, and the fund's expense ratio was 2%, what would your return if you sold the fund at the end of the year be?
Answer #1

Price per Share at beginning = $15

After front end load,

Effective Beginning Price = (1 - 0.0575)(15)

Effective Beginning Price = $14.14

Value at the end of the year = 1.12(14.14)

Value at the end of the year = $15.84

After expense ratio,

Effective value = (0.98)(15.84)

Effective value = 15.52

Rate of Return = (15.52 - 15)/15

Rate of Return = 3.47%